PHYS-2326: University Physics- ame: QUIZ CHAPTER 25: CAPACITORS ) Calculate the equivalent capocitance for the following crcuit 13) Answer: 5.00 μF 6.00 μF 2.00 pF 4.00 u 90.0 V Calculate the equivalent capacitance between the points A and B for the following circuit. All capacitor values are in microfarads. Answer
Answer #1

here,

1)

5 uF and 6 uF are in series , their equivalent capacitance , C1 = 5 * 6 /( 5 +6)

C1 = 2.73 uF

2 uF and 4 uF are in series , their equivalent capacitance , C2 = 2 * 4 /( 2 + 4)

C2 = 1.33 uF

C1 and C2 are in parallel, the equivalent capacitance , Ceq = C1 + C2

C = 4.06 uF

2)

4 uF and 8 uF are in series , their equivalent capacitance , C1 = 4 * 8 /( 4 + 8) = 2.67 uF

C1 , 3 and 4 are in parallel , their equivalent capacitance , C2 = C1 + 3 uF + 4 uF = 9.67 uF

C2 and 6 uF are series , their equivalent capacitance , Ceq = C2 * 6 /( C2 + 6)

Ceq = 9.67 * 6 /( 9.67 + 6)

Ceq = 3.7 uF